Write down the differential equation for a damped harmonic oscillator. What is the basis of representing the damping force in terms of velocity? Show that the average energy of a weakly damped oscillator is given by:
< E > = E0 exp (- 2bt)
1
Expert's answer
2020-02-24T10:47:45-0500

1)


"\sum F=ma"


"F_{damp}=-bv=-b\frac{dx(t)}{dt}"


"F_{el}=-kx(t)"


"-b\frac{dx(t)}{dt}-kx(t)=m\frac{d^2x(t)}{dt^2}"


"\frac{d^2x(t)}{dt^2}+\frac{b}{m}\frac{dx(t)}{dt}+\frac{k}{m}x(t)=0"


So, the differential equation of a damped haronic oscillator


"\frac{d^2x(t)}{dt^2}+2\beta\frac{dx(t)}{dt}+\omega^2_0 x(t)=0" ,


where "\beta=\frac{b}{2m}" and "\omega_0=\sqrt{\frac{k}{m}}"

2)


The forces of resistance and friction are always directed against the direction of the velocity vector and reduce the kinetic energy of the body


3)


"E(t)=KE(t)+PE(t)=\frac{1}{2}m(\frac{dx}{dt})^2+\frac{1}{2}kx^2(t)"


"x(t)=A_0e^{-\beta t}\cos(\omega t+\phi)"


"\frac{dx}{dt}=-A_0e^{-\beta t}[\beta \cos(\omega t+\phi)+\omega \sin(\omega t+\phi)]"


"E(t)=\frac{1}{2}mA^2_0e^{-2\beta t}[(\beta^2+\omega^2_0) \cos^2(\omega t+\phi)+"


"+\omega^2 \sin^2(\omega t+\phi)+\beta \omega \sin2(\omega t+\phi)]"


We have


"\langle E(t) \rangle=\frac{1}{2}mA^2_0e^{-2\beta t}[\frac{\beta^2+\omega^2_0}{2}+\frac{\omega^2}{2}]="


"=\frac{1}{2}mA^2_0e^{-2\beta t}\omega^2_0=E_0e^{-2\beta t}"